OpenStudy (anonymous):

Part A: Solve A = 9 over 2(x + 28) for x. Part B: Determine the value of x when A = 135. Part C: Solve –np – 40 > 10 for n. Show your work.

OpenStudy (anonymous):

\[A= \frac{ 9 }{ 2 } \left( x + 28\right)\]

OpenStudy (anonymous):

@zepdrix can you help me with this?

Vocaloid (vocaloid):

part A: "solve for x" means we want to get x by itself start by multiplying both sides of the equation by 2 to eliminate the denominator

OpenStudy (anonymous):

A2=? (x + 28)

zepdrix (zepdrix):

\[\large\rm A=\frac{9}{2}\left(x+28\right)\]Multiplying both sides by 2 gives you,\[\large\rm 2\cdot A=2\cdot\frac{9}{2}\left(x+28\right)\]Don't think of that 2 as a denominator, think of the operation being performed. It's actually division by 2. So we're multiplying and dividing by 2 on the right side of the equation. Those actions will "cancel out".\[\large\rm 2\cdot A=\cancel{2}\cdot\frac{9}{\cancel{2}}\left(x+28\right)\]

zepdrix (zepdrix):

Which puts you here,\[\large\rm 2A=9(x+28)\]

OpenStudy (anonymous):

divide by 9 and then subtract 28? @zepdrix

OpenStudy (anonymous):

\[\frac{ A2 }{ 9 }-28=x\]

zepdrix (zepdrix):

Ahh sorry had to go make some food +_+

zepdrix (zepdrix):

Looks good!

OpenStudy (anonymous):

2! the answer to B it x = 2!

zepdrix (zepdrix):

yay good job \c:/

OpenStudy (anonymous):

:DD

OpenStudy (anonymous):

-np - 40 > 10 add 40 first or...

zepdrix (zepdrix):

Yah, that seems like a good first step

OpenStudy (anonymous):

-n > 50/p but what about that minus sign?

zepdrix (zepdrix):

I guess you'll want to divide each side by -p, not simply p. Do you remember what happens when you divide a negative across an inequality?

OpenStudy (anonymous):

the inequality flips

zepdrix (zepdrix):

yessss

OpenStudy (anonymous):

n < 50/-p

zepdrix (zepdrix):

Ok great! Let's bring the negative sign out front because \(\large\rm \frac{50}{-p}\) is the same as \(\large\rm \frac{-50}{p}\) which is the same as \(\large\rm -\frac{50}{p}\). It doesn't really matter where the negative goes, it just looks nicer in front like this form.

OpenStudy (anonymous):

ok. do you have time for one more question?

zepdrix (zepdrix):

sure!

OpenStudy (anonymous):

Bruce has a bottle that contains 60% of lemon juice and the rest water. The bottle has 1 liter of water. Part A: Write an equation in one variable that can be used to find the total number of liters of lemon juice and water in the bottle. Define the variable used in the equation. Part B: How many liters of lemon juice are present in the bottle? Show your work. This one is really confusing to me.

zepdrix (zepdrix):

Mmm I remember this one coming up like 6 months ago, confused me also lol gimme a sec XD

OpenStudy (anonymous):

ok :)

OpenStudy (anonymous):

im not so smart

zepdrix (zepdrix):

Oh oh oh oh oh.. ok now I remember why I got confused on this one. Make sure you understand the second sentence. It's NOT saying that bottle contains 1Liter of liquid. It contains 1Liter of water. They're not telling us how much total liquid is in the bottle.

zepdrix (zepdrix):

So umm

OpenStudy (anonymous):

Ok, i see how that can get you!

zepdrix (zepdrix):

Ok if lemon juice accounts for 60% of the liquid in the bottle, what % does that leave for water?

OpenStudy (anonymous):

40%

zepdrix (zepdrix):

Good good good. and I think we can use thattttt :) sec, brain ... not... working lol

OpenStudy (anonymous):

lol

zepdrix (zepdrix):

The `amount` of water is 1Liter. So this 1Liter accounts for 40% of the total Liquid in the bottle. \(\large\rm 1L=.40T\) So how much total liquid is in the bottle? Can you solve this algebraically? (Ignore the L, that's just a unit).

OpenStudy (anonymous):

Darude-Sandstorm

OpenStudy (anonymous):

1.40L

OpenStudy (anonymous):

t

zepdrix (zepdrix):

Woops! The 40% is `multiplying` the Total. So to isoluate your total, T, you'll need to divide.

zepdrix (zepdrix):

isolate*

OpenStudy (anonymous):

2.5

zepdrix (zepdrix):

Good good good. We've determined that the total amount of liquid in the bottle is 2.5Liters. The lemon juice accounts fo 60% of that total. So we can again set this up algebraically:\[\large\rm \ell=.60T\]And since we've solved for the total, T, we can plug it right in,\[\large\rm \ell=.60\cdot 2.5\]And solve for l, the amount of lemon juice in the bottle.

zepdrix (zepdrix):

Do those steps.... kinda make sense? 0_o It's a tricky problem to get through.

OpenStudy (anonymous):

1.5

OpenStudy (anonymous):

1.5=.60*2.5

zepdrix (zepdrix):

Good! :) So there is: 1Liter of water which is 40% of the total liquid 1.5Liter of lemon juice which is 60% of the total liquid 2.5Liters of total liquid which is 1 + 1.5

zepdrix (zepdrix):

Yayyy I think we did it \c:/

OpenStudy (anonymous):

Thank you so much for all your help today! I really understand it better too.

zepdrix (zepdrix):

np \c:/ yay team!